INAF queue observing, 2015 Apr 25-26

Observer: F. Cusano, R. Carini
Telescope Operator: Geno Bechetti
Support Astronomer: Berry Rothberg
Instrument: LBC, LUCI

Summary

At the sunset the sky is mostly cloudy, the wind is too high to open the dome.

Night duration (12\xB0 el twilight) 20:00-04:30 (3:00-11.30 UT) = 8 hrs 30 m

Observing time: 0 hrs

Weather time loss: 8.5 hrs (100%)

Data Summary

At the sunset the sky is mostly cloudy, the wind is too high to open the dome.
3.45 the sky is completely covered
5:10 wind velocity: 25 m/s, humidity 100%

Project_RUN Objects Filter ExpTime Notes
Bias LBC
LUCI Calibration

Details

03:15 Bias for LBC

04:50 Luci Calibrations Dark luci.20150426.0001 test
  • luci.20150426.0002-0006 Piconcelli A6, A5
  • luci.20150426.0007-0011 Piconcelli B1
05:25 LUCI crushed. can not start again. The AS configured LUCI again.

05:45 start again the dark calibrations
  • luci.20150426.0012-0013 test images
  • luci.20150426.0014-0018 Piconcelli B2, Della Ceca D,E,F,G
  • luci.20150426.0019-0023 Rossi E filter HK_spec
  • luci.20150426.0024-0028 Rossi E filter zjspec
  • luci.20150427.0029-0033 standard
09:07 begins flat
  • executeLUCIscript.sh A5_LS_Ks.flat luci.20150427.0034-0043 error in the script
  • executeLUCIscript.sh A5_LS_Ks.flat luci.20150427.0044-0053 not good, the problem is not the script
  • test if the arc script works, but it doesn't
  • the AS and TO do tests to solve the problem

-- %USERSIG{RobertaCarini - 2015-04-26}%

Comments

 
Topic revision: r8 - 29 Apr 2015, RobertaCarini
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback